Quadratic Applications For Falling Bodies
An arrow is launched straight up in the air from the ground. It leaves the bow with an initial velocity of 80ft./sec
a. What is the maximum height it will reach?
b. When will it reach the maximum height?
c. when will it be 50 ft off the ground?
d. how high will it be after 2 seconds?
e. how long will it take for the arrow to come back down and hit the ground?

Answer by Alan3354(69443)

An arrow is launched straight up in the air from the ground. It leaves the bow with an initial velocity of 80ft./sec
a. What is the maximum height it will reach?
b. When will it reach the maximum height?
c. when will it be 50 ft off the ground?
d. how high will it be after 2 seconds?
e. how long will it take for the arrow to come back down and hit the ground?
=========================================================
Do it like this one:
------------
A projectile is launched straight up from the ground level with an initial velocity of 96 ft per second. Its height in feet, "s", after "t" seconds is given by the equation: s=16t^2+96t
------------------
It's -16t^2 + 96t
--------------
a) At what time(s) will it reach a height of 80 feet?
-16t^2 + 96t = 80
Solve for t
-16t^2 + 96t = 80
t^2 - 6t + 5 = 0
(t-5)*(t-1) = 0
t = 1 second (going up)
t = 5 seconts (coming back down)
====================================
b) After how many seconds will it return to the ground?
Same equation, equals 0
solve for t
t = 0 is the time at launch
The 2nd value of t is at impact.